OpenStudy (anonymous):

MEDALSSSS AN FAN!!! Jacob bought seven bottles of water. Each bottle contains one-half liter of water. How many liters did he buy? two and one-seventh three and one-third three and one-half seven and one-half

OpenStudy (mokeira):

\[7 \times \frac{ 1 }{ 2 }\] what do you get?

OpenStudy (anonymous):

\[7\times \frac{ 1 }{ 2}=\frac{ 7 }{ 2 }=3.5=3\frac{ 1 }{ 2}\] Jacob bought three and one-half liters of water.